# Information geometry approach to quantum stochastic thermodynamics

### v2

@@ VI Geometric bound on the von Neumann entropy rate
In this section, we will first demonstrate that the classical uncertainty relation between the geometric uncertainty in the path in state space and the time-averaged rate of information change, established in [35] for classical systems, remains valid for quantum dynamics. We then make use of the decomposition of any QFI into a common metric-independent incoherent contribution and a metric-dependent coherent contribution, revised in section IV. The incoherent part is inherently linked to the system’s entropy dynamics as shown in section V.0.2. This connection enables us to extend a classical bound on the change in the entropy rate, as described in [35], to the quantum domain.
@@
at any time $T$, so that $0\leq T\leq\tau$. The equation above captures the cumulative path deviation from the geodesic, which represents the trajectory of constant speed between the initial and final states. Since $\mathcal{J}-\mathcal{L}^{2}$ vanishes along the geodesic, where the QFI remains constant by definition, any non-zero difference indicates a deviation from this optimal path. Because of the positivity of $\mathcal{L}^{2}$ and $\delta$, one finds that
$\delta\leq\frac{8\mathcal{J}}{T^{2}}.$ (54)
@@
$\frac{\mathcal{I}}{\delta}\geq\frac{1}{2}.$ (55)
The above takes the form of an uncertainty relation, as described in [35], indicating that lower uncertainty in the path comes at the cost of a lower time-averaged rate of information change. We find that this uncertainty relation holds for quantum dynamics in exactly the same way as it does for classical dynamics.
Building on the above uncertainty relation, we decompose the QFI into its incoherent and coherent components, $F_{Q}=F_{Q}^{IC}+F_{Q}^{C}$ and $F_{Q}^{IC}=\mathcal{B}-\ddot{S}$. By performing the integral $\int_{0}^{T}F_{Q}\mathrm{d}t=\int_{0}^{T}\left(\mathcal{B}(t)+F_{Q}^{C}(t)-\ddot{S}(t)\right)\mathrm{d}t$, we find that the difference in the entropy rate between the final and initial states is bounded from above
$\begin{split}\Delta\dot{S}&\leq\mathcal{C}-\frac{\delta T}{2}+\int_{0}^{T}\mathrm{d}tF_{Q}^{C},\\ \end{split}$ (56)
@@
The above bound constitutes the second key result of our work. As in the classical case [35], the change in the entropy rate is constrained by the sum of the integrated difference between the average local information rate and the bulk information rate, as well as the geometric uncertainty. However, an additional non-negative quantum contribution emerges, which we identify as the geometric action associated with the coherent dynamics. It is worth pointing out that all three terms vanish only in stationary states.
Note that if one considers a driving protocol where the initial state is the fixed point of the initial instantaneous generator of the dynamics, such that $\dot{S}(0)=0$, the above serves as a bound on the instantaneous rate of entropy change $\dot{S}(T)$ at any later time $T$, which indicates that the state is out of equilibrium. We can further exploit the relationship between the incoherent contribution to the QFI and the entropic acceleration in Markovian open system dynamics obeying local detailed balance. In cases where the heat flow between the system and its environment vanishes, the resulting bound applies to the entropy production. Conversely, if the entropy production is zero, the bound constrains the entropic flow. It is also worth noting that our bound applies irrespective of the chosen QFI metric. However, it can be shown that $-\frac{\delta T}{2}+\int_{0}^{T}\mathrm{d}tF_{Q}^{C}$ is minimized for the SLD QFI, resulting in the tightest bound.
@@ VII Quantum Mpemba effect
We now apply the information-geometric framework developed in the previous sections to gain insight into the (thermo-)dynamics underlying a quantum Mpemba effect manifest during the thermal relaxation process of a qubit interacting with a thermal bosonic bath, as recently explored in [51]. Historically, the Mpemba effect [75, 76] describes the following phenomenon: consider two identical buckets of liquid. Both are at thermal equilibrium, but with respect to environments at different temperatures, $T_{h}$ and $T_{c}$, with $T_{h}>T_{c}$. Then, these buckets are placed in a third environment with a lower temperature $T_{h}>T_{c}>T^{*}$, and are allowed to thermalise. A Mpemba effect is said to be observed if there exists a time $t_{M}$ after which the temperature of the initially hotter liquid is lower than that of the initially colder liquid at all later times. The effect has stimulated a lot of theoretical and experimental interest [77, 78, 79, 80, 81, 82], however, a (universal) physical origin remains elusive and consensus on whether the effect exists in the first place has not yet been reached [83, 84]. More recently, Mpemba-like effects have been predicted and experimentally observed also in the quantum domain [85, 86, 87, 88, 89], primarily focused on symmetry-restoration dynamics in closed quantum systems.
We address a particular version occurring in the thermalsation behaviour of Markovian open quantum systems, recently put forward by Moroder et al. [51]. The authors used the concept of non-equilibrium free energy $F_{\mathrm{neq}}$ to understand the thermodynamics of the effect. This is defined as
@@
with equilibrium free energy $F_{\mathrm{eq}}=-\beta^{-1}\log Z_{\beta}$, partition function $Z_{\beta}=\mathrm{Tr}\left[\exp(-\beta H)\right]$ and $\mathrm{D}(\hat{\sigma}\|\hat{\tau}_{\beta})$ is the quantum relative entropy between the relaxing state and the thermal fixed point. The authors in [51] defined a Mpemba effect to occur when a $\hat{\rho}_{0}^{\prime}$ with higher initial non-equilibrium free energy $F_{\mathrm{neq}}$ compared to the reference state $\hat{\rho}_{0}$, relaxes to thermal equilibrium faster. In particular, there exists a time $t_{M}$ after which $F_{\mathrm{neq}}(\hat{\rho}^{\prime}(t))>F_{\mathrm{neq}}(\hat{\rho}(t))$ for all $t>t_{M}$, indicating a crossing of non-equilibrium free energies along their respective relaxation paths.
@@
Consider a qubit weakly coupled to a thermal bosonic environment at inverse temperature $\beta$. Interactions between the ground state |$\ket{g}$ and the excited state |$\ket{e}$ separated by energy $\epsilon$ ($\hbar=k_{\mathrm{B}}=1$), with Hamiltonian $\hat{H}=\frac{\epsilon}{2}\hat{\sigma}_{z}$, are facilitated by interactions with a bosonic environment. The transitions occur with rates $\gamma_{+}=\gamma f_{B}(\epsilon,\beta)$ and $\gamma_{-}=\gamma(1+f_{B}(\epsilon,\beta))$, with Bose-Einstein occupation $f_{B}(\epsilon,\beta)=\left[\exp(\beta\epsilon)-1\right]^{-1}$. The GKSL master equation that describes the resulting reduced dynamics of the qubit is given by
@@
The qubit is initialised to an arbitrary mixed state, serving as the reference state $\hat{\rho}_{0}$. It is parameterised as $\hat{\rho}_{0}=\frac{1}{2}\left(1+\vec{r}\cdot\vec{\sigma}\right)$, where $\vec{\sigma}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})^{\mathrm{T}}$ and $\vec{r}=(r_{x},r_{y},r_{z})^{\mathrm{T}}$. We restrict to reference states with coherences in the initial energy eigenbasis so that at least $r_{x}$ or $r_{y}$ are non-zero. The unitarily rotated state, diagonal in the energy eigenbasis, is $\hat{\rho}^{\prime}_{0}=\mathrm{diag}(\lambda_{1},\lambda_{2})$, where $\lambda_{1}$ and $\lambda_{2}$ are the eigenvalues of $\hat{\rho}$, so that $\lambda_{1}\geq\lambda_{2}$. Importantly, the choice of $\hat{\rho}_{0}^{\prime}$ ensures that $F_{\mathrm{neq}}(\hat{\rho}_{0}^{\prime})\geq F_{\mathrm{neq}}(\hat{\rho}_{0})$.
The numerical results, which we will describe shortly, are obtained for the reference state and the rotated state, specified by the vectors $\vec{r}=(-0.41760,-0.60647,0.47879)$ and $\vec{r}^{\prime}=(0,0,0.87836)$, respectively. Additional parameters (with $\hbar=k_{\mathrm{B}}=1$) that determine the set-up are the coupling strength to the bosonic reservoir $\gamma=1$, the temperature $T=10$ and the Hamiltonian level splitting $\omega=5$.
@@
As discussed in section III, $\sqrt{F_{Q}(t)}$ quantifies the speed at which a state moves along its path in state space. At equilibrium, the state is stationary, and therefore, the QFI vanishes. Our findings reveal that at short times, the rotated state (with initially higher non-equilibrium free energy) evolves more rapidly towards equilibrium. In contrast, the coherent reference state reaches the equilibrium state much later, as indicated by the slower decrease in the total QFI, as shown in Fig. 1a). This difference can be understood by examining the incoherent and coherent contributions separately. The coherent contribution to the QFI of the reference state decays exponentially slower than the incoherent contribution, and thus dictates the relaxation timescale. Since the dynamics of populations and coherences are decoupled [90, 91], the decay rates of the incoherent contributions to the QFI are identical for both the reference and rotated states (Fig. 1).
@@
Figure 4: a) Geometric uncertainty $\delta$ about the path of the rotated state $\hat{\rho}^{\prime}$ (red). We compute the geometric uncertainty of the time-evolving reference state $\hat{\rho}$ using the SLD, WY and HM metric (blue). The SLD yields the smallest geometric uncertainty. Deviations from the geodesic connecting the reference state and rotated state to the thermal state, respectively, are accumulated at short times. b) Trade-off between the time-averaged Fisher information and the geometric uncertainty.
@@
Here, we choose the observable of interest to be the Hamiltonian $\hat{H}$, as its changes in the expectation value due to state changes is precisely the heat current, which, for consistency with previously introduced notation, we denote as $\left\lvert\dot{h}(t)\right\rvert$. In Fig. 3, we show the corresponding integrated ratio (black line) in comparison to its bound defined by the statistical distance (blue and red dashed lines). The colored regions represent the areas excluded by these bounds. We first note that since the statistical distance between the initial rotated state $\rho^{\prime}$ and the time-evolved state $\rho^{\prime}(t)$ coincides with the geodesic length between the two at all times, they provide the identical bound (Fig. 3b)). We find further that this bound is saturated at all times, a consequence of the fact that the symmetric logarithmic derivative (see [38]) and the Hamiltonian share the same eigenspaces. Clearly, this is not the case for the initially coherent reference state (Fig. 3a)).
@@
Finally, we examine the trade-off relation between the the time-averaged Fisher information $\mathcal{I}$ and the geometric uncertainty $\delta$. It expresses that lower uncertainty in the path comes at the cost of lower time-averaged rate of information change, as discussed in section VI. Initially, all paths deviate strongly from the geodesic. The bound tells us that the high geometric uncertainty (and thus a high time-variance of the speed of evolution) has to be compensated by a sufficiently large time-averaged speed. We find that this ratio is consistently lower, and thus closer to the bound, during the relaxation of the rotated state at all times, as shown in Fig. 4b).
@@ VIII Conclusion
In our work, we generalize classical results, that connect information geometry to entropic acceleration in systems out of equilibrium, to quantum thermodynamics. Specifically, we demonstrate that for open quantum systems governed by GKSL dynamics, any quantum Fisher information (QFI) is connected to the accelerations of entropy production and entropic flow. To this end, we leverage the fact that any QFI about the time parameter can be decomposed into coherent and incoherent contributions, and we find that the entropic acceleration is expressible solely in terms of the incoherent part. We show that the classical uncertainty relation between the geometric uncertainty in the path in state space and the time-integrated rate of information change is valid also for quantum dynamics. Further, we demonstrate that the classical bound on changes in the von Neumann entropy rate can be extended to the quantum domain through the addition of the non-negative geometric action associated with the coherent dynamics.

### v3

@@ VI Geometric bound on the von Neumann entropy rate
In this section, we will first demonstrate that the classical uncertainty relation between the geometric uncertainty in the path in state space and the time-averaged rate of information change, established in [35] for classical systems, remains valid for quantum dynamics. In fact, we find that the bound, for both classical and quantum systems, can be tightened by a factor of two. We then make use of the decomposition of any QFI into a common metric-independent incoherent contribution and a metric-dependent coherent contribution, revised in section IV. The incoherent part is inherently linked to the system’s entropy dynamics as shown in section V.0.2. This connection enables us to extend a classical bound on the change in the entropy rate, as described in [35], to the quantum domain.
@@
at any time $T$, so that $0\leq T\leq\tau$. Here, $\mathcal{J}$ and $\mathcal{L}$ denote the statistical divergence and statistical length accumulated up to time $T$, respectively. The equation above captures the cumulative path deviation from the geodesic, which represents the trajectory of constant speed between the initial and final states. Since $\mathcal{J}-\mathcal{L}^{2}$ vanishes along the geodesic, where the QFI remains constant by definition, any non-zero difference indicates a deviation from this optimal path. Because of the positivity of $\mathcal{J}$, $\mathcal{L}^{2}$ and $\delta$, one finds that
$\delta\leq\frac{4\mathcal{J}}{T^{2}}.$ (45)
@@
$\frac{\mathcal{I}}{\delta}\geq 1.$ (46)
The above takes the form of an uncertainty relation, indicating that lower uncertainty in the path comes at the cost of a lower time-averaged rate of information change, as argued in [35]. Note, however, that the lower bound derived in [35] is looser than the bound presented here by a factor of $\frac{1}{2}$. This difference arises from the use of the inequality $\mathcal{J}-\mathcal{L}^{2}\leq 2\mathcal{J}$ in their argument, whereas our derivation relies on both $\mathcal{J}-\mathcal{L}^{2}\geq 0$ and $\mathcal{L}^{2}\geq 0$, leading to the tighter bound $\mathcal{J}-\mathcal{L}^{2}\leq\mathcal{J}$. We find that this uncertainty relation holds for both classical and quantum dynamics.
Saturation of the bound can be achieved in the long-time limit during a relaxation process to a steady state, where the QFI with respect to the time parameter vanishes by definition, as the state becomes constant:
[…]
This saturation behavior is discussed and demonstrated in the example of the quantum Mpemba effect in Sec. VII. Building on the above uncertainty relation, we decompose the QFI into its incoherent and coherent components, $F_{Q}=F_{Q}^{IC}+F_{Q}^{C}$ and $F_{Q}^{IC}=\mathcal{B}-\ddot{S}$. By performing the integral $\int_{0}^{T}F_{Q}\mathrm{d}t=\int_{0}^{T}\left(\mathcal{B}(t)+F_{Q}^{C}(t)-\ddot{S}(t)\right)\mathrm{d}t$, we find that the difference in the entropy rate between the final and initial states is bounded from above
$\begin{split}\Delta\dot{S}&\leq\mathcal{C}-T\delta+\int_{0}^{T}\mathrm{d}tF_{Q}^{C},\\ \end{split}$ (48)
@@
The above bound constitutes the second key result of our work. Similar to the classical case [35], the change in the entropy rate is constrained by the sum of the integrated difference between the average local information rate and the bulk information rate, as well as the geometric uncertainty. However, an additional non-negative quantum contribution emerges, which we identify as the geometric action associated with the coherent dynamics. As briefly described in Sec. IV, the action is associated to the ‘energy’ of the path due to its similarity with the kinetic energy integral in classical mechanics. It is worth pointing out that the first two terms vanish only in stationary states [35], whereas the last tern vanishes in stationary states or whenever the dynamics is purely incoherent. Therefore, their sum vanishes only in stationary states. Note that if one considers a driving protocol where the initial state is the fixed point of the initial instantaneous generator of the dynamics, such that $\dot{S}(0)=0$, the above serves as a bound on the instantaneous rate of entropy change $\dot{S}(T)$ at any later time $T$, which indicates that the state is out of equilibrium. We can further exploit the relationship between the incoherent contribution to the QFI and the entropic acceleration in Markovian open system dynamics obeying local detailed balance. In cases where the heat flow between the system and its environment vanishes, the resulting bound applies to the entropy production. Conversely, if the entropy production is zero, the bound constrains the entropic flow.
It is also worth noting that our bound applies irrespective of the chosen QFI metric. However, it can be shown that since $-\delta T+\int_{0}^{T}\mathrm{d}tF_{Q}^{C}=4\mathcal{L}_{f}^{2}/T-\int_{0}^{T}\mathrm{d}tF_{Q}^{IC}$, it is minimized for the SLD QFI, resulting in the tightest bound.
@@ VII Quantum Mpemba effect
We now apply the information-geometric framework developed in the previous sections to analyze the (thermo-)dynamics underlying a quantum Mpemba effect. This effect is manifest during the thermal relaxation of a qubit interacting with a thermal bosonic bath and was recently explored in [52]. Historically, the Mpemba effect [79, 80] describes the following phenomenon: consider two identical buckets of liquid. Both are at thermal equilibrium, but with respect to environments at different temperatures, $T_{h}$ and $T_{c}$, with $T_{h}>T_{c}$. Then, these buckets are placed in a third environment with a lower temperature $T_{h}>T_{c}>T^{*}$, and are allowed to thermalise. A Mpemba effect is said to be observed if there exists a time $t_{M}$ after which the temperature of the initially hotter liquid is lower than that of the initially colder liquid at all later times. The effect has stimulated a lot of theoretical and experimental interest [81, 82, 83, 84, 85, 86, 87], however, a (universal) physical origin remains elusive and consensus on whether the effect exists in the first place has not yet been reached [88, 89]. More recently, Mpemba-like effects have been predicted and experimentally observed also in the quantum domain [90, 91, 92, 93, 94], primarily focused on symmetry-restoration dynamics in closed quantum systems.
We address a particular version occurring in the thermalisation behaviour of Markovian open quantum systems, recently put forward by Moroder et al. [52]. The authors used the concept of non-equilibrium free energy $F_{\mathrm{neq}}$ to understand the thermodynamics of the effect. This is defined as
@@
with equilibrium free energy $F_{\mathrm{eq}}=-\beta^{-1}\log Z_{\beta}$, partition function $Z_{\beta}=\mathrm{Tr}\left[\exp(-\beta\hat{H})\right]$ and $\mathrm{D}(\hat{\sigma}\|\hat{\tau}_{\beta})$ is the quantum relative entropy between the relaxing state and the thermal fixed point $\hat{\tau}_{\beta}=\exp(-\beta\hat{H})/Z_{\beta}$, with respect to the system Hamiltonian $\hat{H}$. The authors in [52] defined a Mpemba effect to occur when a $\hat{\rho}_{0}^{\prime}$ with higher initial non-equilibrium free energy $F_{\mathrm{neq}}$ compared to the reference state $\hat{\rho}_{0}$, relaxes to thermal equilibrium faster. In particular, there exists a time $t_{M}$ after which $F_{\mathrm{neq}}(\hat{\rho}(t))>F_{\mathrm{neq}}(\hat{\rho}^{\prime}(t))$ for all $t>t_{M}$, indicating a crossing of non-equilibrium free energies along their respective relaxation paths.
@@
Consider a qubit weakly coupled to a thermal bosonic environment at inverse temperature $\beta$. Interactions between the ground state |$\ket{g}$ and the excited state |$\ket{e}$ separated by energy $\epsilon$ ($\hbar=k_{\mathrm{B}}=1$), with Hamiltonian $\hat{H}=\frac{\epsilon}{2}\hat{\sigma}_{z}$, are facilitated by interactions with a bosonic environment. The transitions occur with rates $\gamma_{+}=\gamma f_{B}(\epsilon,\beta)$ and $\gamma_{-}=\gamma(1+f_{B}(\epsilon,\beta))$, with Bose-Einstein occupation $f_{B}(\epsilon,\beta)=\left[\exp(\beta\epsilon)-1\right]^{-1}$ and coupling strength $\gamma$. The GKSL master equation that describes the resulting reduced dynamics of the qubit is given by
@@
The qubit is initialised to an arbitrary mixed state, serving as the reference state $\hat{\rho}_{0}$. It is parameterised as $\hat{\rho}_{0}=\frac{1}{2}\left(1+\vec{r}\cdot\vec{\sigma}\right)$, where $\vec{\sigma}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})^{\mathrm{T}}$ and $\vec{r}=(r_{x},r_{y},r_{z})^{\mathrm{T}}$. We restrict to reference initial states with coherences in the energy eigenbasis so that at least $r_{x}$ or $r_{y}$ are non-zero. A unitarily rotated state, diagonal in the energy eigenbasis, is $\hat{\rho}^{\prime}_{0}=\mathrm{diag}(\lambda_{1},\lambda_{2})$, where $\lambda_{1}$ and $\lambda_{2}$ are the eigenvalues of $\hat{\rho}$, so that $\lambda_{1}\geq\lambda_{2}$. Importantly, this choice of $\hat{\rho}_{0}^{\prime}$ ensures that $F_{\mathrm{neq}}(\hat{\rho}_{0}^{\prime})\geq F_{\mathrm{neq}}(\hat{\rho}_{0})$.
The numerical results, which we will describe shortly, are obtained for the reference state and the rotated state, specified by the vectors $\vec{r}=(-0.41760,-0.60647,0.47879)$ and $\vec{r}^{\prime}=(0,0,0.87836)$, respectively. Additional parameters (with $\hbar=k_{\mathrm{B}}=1$) that determine the set-up are the coupling strength to the bosonic reservoir $\gamma=1$, the temperature $T=10$ and the Hamiltonian level splitting $\epsilon=5$.
@@
As discussed in section III, $\sqrt{F_{Q}(t)}$ quantifies the speed at which a state moves along its path in state space. At equilibrium, the state is stationary, and therefore, the QFI vanishes. Our findings reveal that at short times, the rotated state (with initially higher non-equilibrium free energy) evolves more rapidly towards equilibrium. In contrast, the coherent reference state reaches the equilibrium state much later, as indicated by the slower decrease in the total QFI, as shown in Fig. 1a). This difference can be understood by examining the incoherent and coherent contributions separately. The coherent contribution to the QFI of the reference state, due to Hamiltonian dynamics, decays exponentially slower than the incoherent contribution stemming from the dissipative dynamics, and thus dictates the relaxation timescale. Since the dynamics of populations and coherences are decoupled [95, 96], the decay rates of the incoherent contributions to the QFI are identical for both the reference and rotated states (Fig. 1).
@@
Figure 4: a) Geometric uncertainty $\delta$ about the path of the rotated state $\hat{\rho}^{\prime}$ (red). We compute the geometric uncertainty of the time-evolving reference state $\hat{\rho}$ using the SLD, WY and HM metric (blue). The SLD yields the smallest geometric uncertainty. Deviations from the geodesic connecting the reference state and rotated state to the thermal state, respectively, are accumulated at short times. b) Trade-off between the time-averaged QFI and the geometric uncertainty.
@@
Here, we choose the observable of interest to be the Hamiltonian $\hat{H}$, as its changes in the expectation value due to state changes is precisely the heat current, which, for consistency with previously introduced notation, we denote as $\left\lvert\dot{h}(t)\right\rvert$. In Fig. 3, we show the corresponding integrated ratio (black line) in comparison to its bound defined by the statistical distance (blue and red dashed lines). The colored regions represent the areas excluded by these bounds. We first note that since the statistical distance between the initial rotated state $\rho^{\prime}$ and the time-evolved state $\rho^{\prime}(t)$ coincides with the geodesic length between the two at all times, they provide the identical bound (Fig. 3b)). We find further that this bound is saturated at all times, a consequence of the fact that along the incoherent trajectory, the covariance between the logarithmic derivative $\hat{L}_{\text{SLD}}$ and the Hamiltonian $\hat{H}$ factorises into the the product of their respective standard deviation with respect to $\hat{\rho}^{\prime}_{t}$, so that $|\dot{h}(t)|=\text{cov}_{\hat{\rho}^{\prime}_{t}}(\hat{H},\hat{L}_{\text{SLD}})=\Delta_{\hat{\rho}^{\prime}_{t}}\hat{H}\Delta_{\hat{\rho}^{\prime}_{t}}\hat{L}_{\text{SLD}}=\Delta_{\hat{\rho}^{\prime}_{t}}\hat{H}\sqrt{F_{Q}(t)}$, as discussed in detail in [38]. Clearly, this is not the case for the initially coherent reference state (Fig. 3a)).
@@
Finally, we examine the trade-off relation between the the time-averaged Fisher information $\mathcal{I}$ and the geometric uncertainty $\delta$. It expresses that lower uncertainty in the path comes at the cost of lower time-averaged rate of information change, as discussed in section VI. Initially, all paths deviate strongly from the geodesic. The bound tells us that the high geometric uncertainty (and thus a high time-variance of the speed of evolution) has to be compensated by a sufficiently large time-averaged speed. We find that this ratio is consistently lower, and thus closer to the bound, during the relaxation of the rotated state at all times, as shown in Fig. 4b). At long times, the bound approaches saturation because the state nears the thermal steady state of the dynamics, where $F_{Q}$ vanishes due to the state’s time invariance, as discussed in Sec. VI.
@@ VIII Conclusion
In our work, we present a fully quantum mechanical derivation that connects information-geometric quantities, defined via the quantum Fisher information (QFI) with respect to the time parameter, to quantum thermodynamics within the framework of stochastic processes. This enables us to extend classical results on entropy dynamics in nonequilibrium systems to the quantum domain. Specifically, we demonstrate that for open quantum systems governed by GKSL dynamics, any QFI is connected to the accelerations of entropy production and entropic flow. To this end, we leverage the fact that any QFI about the time parameter can be decomposed into coherent and incoherent contributions, and we find that the entropic acceleration is expressible solely in terms of the incoherent part. We tighten the classical uncertainty relation between the geometric uncertainty in the path in state space and the time-integrated rate of information change, and show that it is valid also for quantum dynamics. Further, we demonstrate that the classical bound on changes in the von Neumann entropy rate can be extended to the quantum domain through the addition of the non-negative geometric action associated with the coherent dynamics. While this bound a priori holds for any QFI, we find that it is tightest for the SLD QFI.